OpenStudy (aaronandyson):

5tan - 2/6 + tan where tan is 2/3

OpenStudy (adamaero):

What do you mean tan = 2/3 ?

OpenStudy (aaronandyson):

yes

OpenStudy (aaronandyson):

tan = 2/3

OpenStudy (adamaero):

I'm guessing you mean tan(x) = 2/3 and not that tan is a variable?

OpenStudy (aaronandyson):

Take it as tan(a)

OpenStudy (adamaero):

Just repost the question exactly as it's posed. It's just confusing of you make shortcuts in notation.